OpenStudy (anonymous):

A right triangle’s legs are 3 and 4 meters long. What is the measure of the angle adjacent to the 4-meter leg to the nearest tenth of a degree?

OpenStudy (anonymous):

* 35.9 *36 *36.9 *37.5

jhonyy9 (jhonyy9):

ok so there is a right triangle what you know from the text of this your exercise yes ?

jhonyy9 (jhonyy9):

do you know the pythagorean numbers ?

jhonyy9 (jhonyy9):

what is the theorem of Pythagoras ?

OpenStudy (anonymous):

yes there is a right angle. but that is the only thing about this that i understand

jhonyy9 (jhonyy9):

so than you dont know the theorem of Pythagoras ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

so the theorem of Pythagora say that in case of a right triangle the square of hypothenuse is equal with the sum of the square of the shortly sides let a and b the shortly sides and c is the hypothenuse so than there is c^2 = a^2 + b^2 ok ?

OpenStudy (anonymous):

yeah i get that part.

jhonyy9 (jhonyy9):



OpenStudy (anonymous):

25?

jhonyy9 (jhonyy9):

25 is the square of x but you need calculi the x

OpenStudy (anonymous):

3+4=7?

jhonyy9 (jhonyy9):

no check it above how i have wrote there the Pythagorean theorem

OpenStudy (anonymous):

how do i figure A?

jhonyy9 (jhonyy9):

x^2 = a^2 + b^2

jhonyy9 (jhonyy9):

x^2 = 3^2 +4^2

jhonyy9 (jhonyy9):

x^2 = 9+16 =25 x^2 =25

OpenStudy (anonymous):

\[3^{2}+4^{2}=25\]

jhonyy9 (jhonyy9):

so than x = ?

OpenStudy (anonymous):

12.5?

jhonyy9 (jhonyy9):

x^2 = 25 x = sqrt25 = ?

jhonyy9 (jhonyy9):

how many is sqrt25 ?

OpenStudy (anonymous):

5

jhonyy9 (jhonyy9):

yes right

OpenStudy (anonymous):

but what is the measure of the angle adjacent to the 4-meter leg to the nearest tenth of a degree

jhonyy9 (jhonyy9):

so this will be angle C on my draw yes ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

do you know how many will be sin C = ? what is the sinus of an angle ?

jhonyy9 (jhonyy9):

sorry this was wrong so the sinus of an angle is opposite side divide hypothenuse can you rewrite and calculi it using these values of sides ?

jhonyy9 (jhonyy9):

sin C = 3/5

jhonyy9 (jhonyy9):

sin C = 0,6

OpenStudy (anonymous):

you lost me

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

i am lost on sin cas and tan

jhonyy9 (jhonyy9):

than you need learning again some , trigonometry from your math book yes ?